This is our first weekly report after the 6 week accelerator site shut
down. Highlights for the week include initial checkout of the FEL driver
accelerator with electron beam in preparation for a one month run starting on Monday, July 12.
During the month of June the project hosted a two day workshop
(on June 17-18th) for the Laser Processing Consortium, the Industrial
Advisory Board and prospective FEL users. The workshop was attended by 80 participants (41 from universities, 23 from National Labs, 13 from industry and 3 from government). The agenda included : a status report on the FEL performance; procedures for preparing FEL users tests; reports from the first three user tests performed during the May run, reports on FEL user proposals that are in review or preparation. At the LPC workshop we were pleased to announce that four user groups had received recent announcements of grants for FEL experiments. Congratulations to L.Feldman and N.Tolk at Vanderbilt (who received a DOE-BES grant); to A.Reilly at William & Mary (NSF grant); E. Gillman, NSU and G. Neil, Jlab (ONR grant); and J. Clarke, Northrop Grumman (NASA grant).
Preparations for the FEL Upgrade project continued during the last month with our work on a draft technical proposal, draft requirements document and continued refinement of the cost estimate
FEL Installation/Maintenance Activities
Hardware changes made during the shut down period include installing: a new cathode ball and wafer in the gun (after which resurrection of the gun proceeded smoothly), red filters on the CCD cameras used for quadrupole/viewer emittance measurements (they should cut down the background from ghost pulses and provide more accurate data), and three synchrotron light monitors, one in the first optical chicane and one in each recirculation arc, to be used to monitor stability of the electron beam. In addition, much work was done in further configuring User Labs 1, 3, and 4, including addition of a remote-controlled "turnstile" that will be used to insert new samples rapidly into the laser beamline. A number of software changes were also made during the down, but the only one of direct interest to an operator is the addition of an FEL RF Alarm Page.
FEL Commissioning Activities
As of this week, the gun is up and running; its voltage is being kept at a conservative 320 kV to keep the dark current down, to take stress off of its performance/availability, and to enable immediate use of the machine settings from the last run. The reserve is there (up to ~340-350 kV) should we ever wish to push up its voltage, e.g., in an effort to improve beam quality. At this writing (1330 Friday, 9 July 99), we are in process of setting up the injector with pulsed beam. Plans are to run Day Shift today to complete the injector setup and run some cw beam. The machine would then be ready for next week's program. At present, the one persistent difficulty is frequent trips of the waveguide IR interlocks to the cryomodule cavities. During the down, screens were installed in each waveguide to block rf radiation thought to be causing the trips, but preliminary indications are that they are not helpful. We are reconsidering how to rectify the problem.
A brief summary of plans for the next two weeks follows: Because the new sapphire ( 3 micron) mirrors are not yet ready, plans are to install them Tuesday of next week. Prior to that, on Monday, plans are to push the average current (without lasing) to ~5 mA, and then after the mirrors are in place, we will try again for "1.0 kW" lasing. Coherent synchrotron radiation studies begin next week and continue through the first three weeks of Owl Shifts; detailed plans were developed during the down. Beam- breakup studies will begin on or about 19 July 99 and will occupy two weeks of Day Shifts. Lasing studies (laser diagnostics and high-priority user tests) will also begin 19 July and occupy two weeks of Swing Shifts. As presently envisioned, the fourth week of the run will be dedicated exclusively to injector studies. Of course, depending on how things progress, there could be adjustments to the program later, but ideally things will proceed in the manner described.
Courtesy of D. Douglas, the latest thinking on the beam transport for
the Upgrade, which is probably the most difficult aspect of the new
machine, is to use new Bates bends for recirculation, thereby cutting both cost and risk by building on what we know from the present
machine. The biggest uncertainty is in achieving a full 10% acceptance; third-order effects enter into consideration, and existing codes are
inaccurate beyond second order. Top-level performance goals for the Upgrade are now developed, and we continue to think of ways to reduce cost and risk without sacrificing the performance goals.